# Uniform irreducibility of Galois action on the $\ell$-primary part of Abelian $3$-folds of Picard type

### v1

@@ 3.1. A lemma on surfaces with isolated singularities
[…]
@@
the pullback homomorphism on cohomology. By [6, Thm.3.2.5(iii)] we know that $\widetilde{j}^{\ast}$ is a homomorphism of mixed Hodge structure, with $\operatorname{H}^{1}(\widetilde{X},\mathbf{Q})$ being pure of weight $1$.
@@
by [9, Thm.5.4.12] $j^{\ast}:\mathrm{IH}^{1}(X,\mathbf{Q})\to\mathrm{IH}^{1}(U,\mathbf{Q})$ is an isomorphism, while $\widetilde{j}^{\ast}$ is injective. Moreover by Cor.5.4.11 and Prop.5.4.4 in loc.cit. $\phi^{\ast}:\mathrm{IH}^{1}(X,\mathbf{Q})\to\mathrm{IH}^{1}(\widetilde{X},\mathbf{Q})=\operatorname{H}^{1}(\widetilde{X},\mathbf{Q})$ is an embedding, while $\mathrm{IH}^{1}(U,\mathbf{Q})=\operatorname{H}^{1}(U,\mathbf{Q})$. This is summarized in the following commutative diagram:
@@
3.2. A formula for the irregularity
Let $z\mapsto\bar{z}$ be the non-trivial automorphism of $M/\mathbf{Q}$. Put $M^{1}=\{z\in M^{\times}\mid z\bar{z}=1\}$, which we will view as an algebraic torus over $\mathbf{Q}$ and denote by $\mathbf{A}_{M}^{1}$ its adelic points.
For $\widetilde{K}$ an open compact subgroup of $\widetilde{G}(\mathbf{A}_{f})$ we recall the Shimura variety of Picard type defined as the adelic quotient
@@ 3.2. A formula for the irregularity
[…]
(11) $\Gamma=\widetilde{G}(\mathbf{Q})\cap\widetilde{K}\widetilde{G}(\mathbf{R}),$
@@
(12) $q(Y_{\Gamma}^{\ast})=\sum_{(\lambda,\nu)\in\Xi/\widehat{\pi_{0}(Y_{K})}}\sum_{\pi_{f}\in\Pi_{f}(\lambda,\nu)}\dim(\pi_{f}^{K})\frac{1+W(\lambda\nu_{M})(-1)^{s(\pi_{f})}}{2},\text{ where }$
  • $\Xi$ is the set of pairs $(\lambda,\nu)$ of a unitary Hecke character $\lambda$ of $M$ whose restriction to $\mathbf{Q}$ is $\left(\tfrac{\cdot}{D}\right)$, and of a unitary character $\nu$ of $\mathbf{A}_{M}^{1}/M^{1}$, such that
@@
  • $\Pi_{f}(\lambda,\nu)$ is the finite part of a global Arthur packet for $G$ (see §1.2),
  • $W(\lambda\nu_{M})\in\{\pm 1\}$ is the global root number, where $\nu_{M}(z)=\nu(\bar{z}/z)$ for $z\in\mathbf{A}_{M}^{\times}$,
@@
  • $\mu\in\widehat{\pi_{0}(Y_{K})}$ acts freely on $\Xi$ by sending $(\lambda,\nu)$ to $(\lambda\mu_{M}^{-1},\nu\mu)$.
@@
We show that $Y_{\Gamma}^{\ast}$ admits only isolated singularities and we first observe that the complement of $Y_{\Gamma}$ in $Y_{\Gamma}^{\ast}$ consists of finitely many points, the cusps. A singular point of $Y_{\Gamma}$ which is not an elliptic point, is necessarily a fixed point of a single complex reflexion (an order $2$ element in $\Gamma$ fixing a hyperbolic plane). Although the universal cover $\mathcal{H}\to Y_{\Gamma}$ is not étale at such a point, this is still a smooth point on the quotient, as locally in the analytic geometry the complex reflexion sends $(\tau,z)$ to $(\tau,-z)$ (see [16, §4.5] for more details and additional material). Thus there exists a smooth open $U_{K}$ of the normal projective surface $Y_{K}^{\ast}$ whose complement consists of finitely many closed points. Lemma 3.1 applied component-wise to $U_{K}$ yields $q(Y_{K}^{\ast})=\dim\operatorname{H}^{0}(U_{K},\Omega^{1}_{U_{K}})$. Let $K^{\prime}$ be any normal finite index torsion free subgroup of $K$, e.g. the intersection with the principal congruence subgroup of level $3$ (see [10, Lem. 1.4]). By Koecher’s Principle, as $Y_{K^{\prime}}\backslash U_{K^{\prime}}$ has codimension at least $2$ in $Y_{K^{\prime}}$, we have
@@
where $U_{K^{\prime}}$ is the inverse image of $U_{K}$ under the natural projection $Y_{K^{\prime}}\to Y_{K}$. Taking invariants by the finite group $K/K^{\prime}$ in Rogawski’s formula [10, (14)] for $q(Y_{K^{\prime}}^{\ast})=\dim\operatorname{H}^{0}(Y_{K^{\prime}},\Omega^{1}_{Y_{K^{\prime}}})$ yields
$\dim\operatorname{H}^{0}(U_{K},\Omega^{1}_{U_{K}})=\sum_{(\lambda,\nu)\in\Xi}\sum_{\pi_{f}\in\Pi(\lambda_{f},\nu_{f})}\dim(\pi_{f}^{K})\frac{1+W(\lambda\nu_{M})(-1)^{s(\pi_{f})}}{2}.$
One should note a misprint in loc. cit. where one should read $(1+W(\lambda\nu_{M})(-1)^{s(\pi_{f})})$ instead of $(W(\lambda\nu_{M})+(-1)^{s(\pi_{f})})$. The presence of this root number translates the fact that for $\pi_{f}\in\Pi(\lambda_{f},\nu_{f})$ and $\pi_{\infty}$ the unique non-tempered holomorphic representation in the local Arthur packet $\Pi(\lambda_{\infty},\nu_{\infty})$, $\pi=\pi_{f}\otimes\pi_{\infty}$ is automorphic if, and only if, $W(\lambda\nu_{M})=(-1)^{s(\pi_{f})}$. Both $\dim(\pi_{f}^{K})$ and $1+W(\lambda\nu_{M})(-1)^{s(\pi_{f})}$ being preserved by the action of $\widehat{\pi_{0}(Y_{K})}$, one deduces the desired formula for $q(Y_{\Gamma}^{\ast})$ as in [10, (15)]. ∎
@@ 3.3. Twists of canonical characters and root numbers
Hecke characters $(\lambda,\nu)\in\Xi$ whose local components at each finite place have ‘minimal’ ramification are intimately related to the canonical characters studied by Gross and Rohrlich. They play a pivotal role in our production of automorphic forms contributing to the irregularity of the Picard modular surfaces of low level. We will now briefly recall some of their properties under the assumption that $D>3$ is odd. Consider the character $\lambda_{\infty}(z)=\bar{z}\cdot|z|^{-1}$ of $M_{\infty}^{\times}\simeq\mathbf{C}^{\times}$ and let $\lambda_{f}:\widehat{\mathcal{O}}_{M}^{\times}\to\mathbf{C}^{\times}$ be a continuous character whose restriction to $\mathcal{O}_{M,p}^{\times}$ is given by the unique quadratic character
@@
for all $p$ dividing $D$, and is trivial otherwise. As $\left(\tfrac{-1}{D}\right)=-1$, it follows that $\lambda_{\infty}$ and $\lambda_{f}$ agree on $\mathcal{O}_{M}^{\times}=\{\pm 1\}$. The finiteness of the class group $\mathscr{C}\!\ell_{M}=\mathbf{A}_{M}^{\times}/M^{\times}\widehat{\mathcal{O}}_{M}^{\times}M_{\infty}^{\times}$ guarantees that the resulting character of $M^{\times}\widehat{\mathcal{O}}_{M}^{\times}M_{\infty}^{\times}$ extends to a character $\lambda$ of $\mathbf{A}_{M}^{\times}$ and clearly two such extensions must differ by a class character. As by construction the restriction of $\lambda_{f}$ to $\widehat{\mathbf{Z}}^{\times}$ agrees with the quadratic Dirichlet character $\left(\tfrac{\cdot}{D}\right)$ viewed as a character of $\mathbf{A}^{\times}/\mathbf{Q}^{\times}\mathrm{Nm}(\mathbf{A}_{M}^{\times})=\operatorname{Gal}(M/\mathbf{Q})$, and $\mathbf{A}^{\times}=\mathbf{Q}^{\times}\widehat{\mathbf{Z}}^{\times}\mathbf{R}_{+}^{\times}$, it follows that the restriction of $\lambda$ to $\mathbf{A}^{\times}$ equals $\left(\tfrac{\cdot}{D}\right)$ (i.e. $\lambda$ is conjugate-symplectic). Such a character is called a canonical character and we will denote it by $\lambda_{c}$, remembering that it is only unique up to a multiplication by a character of $\mathscr{C}\!\ell_{M}$. The root number $W(\lambda_{c}^{3})=-W(\lambda_{c})=\left(\tfrac{-2}{D}\right)$ is $1$ if, and only if, $D\equiv 3\pmod{8}$ (see [27]).
@@
Assume henceforth that $\det(K)=\widehat{\mathcal{O}}_{M}^{1}$, so that $\pi_{0}(Y_{K})=\mathscr{C}\!\ell_{M}^{1}$. Assuming further that $\widehat{\mathcal{O}}_{M}^{1}$ embeds centrally into $K$, the central character $\omega=\nu\cdot\lambda_{|M^{1}}$ of any $\pi$ contributing to (12) has to be everywhere unramified, i.e.,
@@
(13) $q(Y_{\Gamma}^{\ast})=\frac{1}{|\mathscr{C}\!\ell_{M}^{1}|}\sum_{\chi\in\Xi^{1},\omega\in\widehat{\mathscr{C}\!\ell_{M}^{1}}}\sum_{\pi_{f}\in\Pi_{f}(\lambda_{c}\chi_{M},\lambda_{c|M^{1}}^{-1}\chi^{2}\omega)}\dim(\pi_{f}^{K})\frac{1+W(\lambda_{c}^{3}\chi_{M}^{3})(-1)^{s(\pi_{f})}}{2},$
@@
where $\Xi^{1}$ denotes the set of finite order characters of $\mathbf{A}_{M}^{1}/M^{1}$ (see [27, (3)] for the fact that multiplication by a class character does not change the root number).
@@
If $3$ does not divide $|\mathscr{C}\!\ell_{M}^{1}|$, then the action of $\mu\in\widehat{\mathscr{C}\!\ell_{M}^{1}}=\mathbf{A}_{M}^{1}/M^{1}\widehat{\mathcal{O}}_{M}^{1}M_{\infty}^{1}$ sending $(\chi,\omega)$ to $(\chi\mu^{-1},\omega\mu^{3})$ allows to twist out the central character and obtain the simpler formula:
(14) $q(Y_{\Gamma}^{\ast})=\sum_{\chi\in\Xi^{1}}\sum_{\pi_{f}\in\Pi_{f}(\lambda_{c}\chi_{M})}\dim(\pi_{f}^{K})\frac{1+W(\lambda_{c}^{3}\chi_{M}^{3})(-1)^{s(\pi_{f})}}{2},$
where $\Pi_{f}(\lambda)$ is a short-hand for $\Pi_{f}(\lambda,\lambda_{|M^{1}}^{-1})$. Proving this formula in general amounts to showing that $\dim(\pi_{f}^{K})$ remains unchanged when multiplying $\lambda$ or $\nu$ by class characters, which we will later check in all cases of interest.
@@
Successfully applying (14) requires to understand how root numbers behave under twisting. As we are interested in creating irregularity at level $\Gamma^{\prime\prime}_{0}(\ell^{r})$, we focus on characters $\chi$ which are only ramified at the fixed inert prime $\ell$.
@@
Assuming that $\chi_{M}^{3}$ has Artin conductor $\ell^{a}$, we have
$W(\lambda_{c}^{3}\chi_{M}^{3})=(-1)^{a}\chi_{\ell}(-1)W(\lambda_{c}^{3}).$
@@
Using the factorization of root numbers $W=\prod_{v}W_{v}$, it suffices to prove that
@@
(16) $W_{v}(\lambda_{c}^{3}\chi_{M}^{3})$ $=W_{v}(\lambda_{c}^{3}\chi_{M}^{3}),\,\text{ for all } v\neq\ell,$
where the local factors are defined using the standard additive character $\psi_{M}=\psi_{\mathbf{Q}}\circ\operatorname{Tr}_{M/\mathbf{Q}}$. Applying [28, Prop. 3] to both $\lambda_{c,\ell}^{3}$ and $\lambda_{c,\ell}^{3}\chi_{M,\ell}^{3}$ yields (15). As $\chi_{M,\infty}=1$, it suffices to check (16) for $v$ finite. Moreover, the characters $\lambda_{c,v}^{3}$ and $\chi_{M,v}$ are unramified for $v\nmid\ell D$, hence both sides of (16) are $1$. Finally, for $v$ dividing $D$, $\chi_{M,v}$ is unramified, $\lambda_{c,v}^{3}$ is tamely ramified, while the additive character $\psi_{E}$ has conductor $1$, implying by [7, (5.5.1)] that $W_{v}(\lambda_{c}^{3}\chi_{M}^{3})$ and $W_{v}(\lambda_{c}^{3})$ differ by $\chi_{M,v}^{3}(\xi^{1+1\cdot 1})=(\chi_{v}(-1))^{2}=1$. ∎

### v2

@@ 3.1. A lemma on surfaces with isolated singularities
We denote by $q(X)$ the irregularity of a projective algebraic surface $X$, given by the dimension of $\operatorname{H}^{1}(X,\mathcal{O}_{X})$. If $X$ is smooth and projective over $\mathbf{C}$, then $q(X)=\dim\operatorname{H}^{0}(X,\Omega^{1}_{X})$.
Let $X$ be a projective irreducible algebraic surface over $\mathbf{C}$ with isolated singularities, i.e., such that there exists a smooth open $j:U\hookrightarrow X$ whose complement $Z=X{\smallsetminus}U$ consists of finitely many closed points. There exists a smooth resolution
@@
the pullback homomorphism on Betti cohomology. By [6, Thm. 3.2.5(iii)] we know that $\widetilde{j}^{\ast}$ is a homomorphism of mixed Hodge structure, with $\operatorname{H}^{1}(\widetilde{X},\mathbf{Q})$ being pure of weight $1$.
@@
by [10, Thm. 5.4.12] $j^{\ast}:\mathrm{IH}^{1}(X,\mathbf{Q})\to\mathrm{IH}^{1}(U,\mathbf{Q})$ is an isomorphism, while $\widetilde{j}^{\ast}$ is injective. Moreover by Corollary 5.4.11 and Proposition 5.4.4 in loc.cit. $\phi^{\ast}:\mathrm{IH}^{1}(X,\mathbf{Q})\to\mathrm{IH}^{1}(\widetilde{X},\mathbf{Q})=\operatorname{H}^{1}(\widetilde{X},\mathbf{Q})$ is an embedding, while $\mathrm{IH}^{1}(U,\mathbf{Q})=\operatorname{H}^{1}(U,\mathbf{Q})$. This is summarized in the following commutative diagram:
@@
3.2. Irregularity of Picard modular surfaces
Let $z\mapsto\bar{z}$ be the non-trivial automorphism of $M/\mathbf{Q}$. Put $M^{1}=\{z\in M^{\times}\mid z\bar{z}=1\}$ viewed as an algebraic torus over $\mathbf{Q}$ and denote by $\mathbf{A}_{M}^{1}$ its adelic points.
@@ 3.2. Irregularity of Picard modular surfaces
$Y_{K}(\mathbf{C})=G(\mathbf{Q})\backslash\left(\mathcal{H}\times G(\mathbf{A}_{f})/K\right),$
@@
(17) $q(Y_{\Gamma}^{\ast})=\sum_{(\lambda,\nu)\in\Xi/\widehat{\pi_{0}(Y_{K})}}\sum_{\pi_{f}\in\Pi_{f}(\lambda,\nu)}\dim(\pi_{f}^{K})\frac{1+W(\lambda\nu^{-1}_{M})(-1)^{s(\pi_{f})}}{2},\text{ where }$
  • $\Xi$ is the set of pairs $(\lambda,\nu)$ of a Hecke character $\lambda$ of $M$ whose restriction to $\mathbf{Q}$ is $\left(\tfrac{\cdot}{D}\right)$ and of a character $\nu$ of $\mathbf{A}_{M}^{1}/M^{1}$, such that
@@
  • $\Pi_{f}(\lambda,\nu)$ is the finite part of the global Arthur packet from §1.2,
  • $W(\lambda\nu_{M}^{-1})\in\{\pm 1\}$ is the global root number, where $\nu_{M}(z)=\nu(z/\bar{z})$ for $z\in\mathbf{A}_{M}^{\times}$,
@@
  • $\mu\in\widehat{\pi_{0}(Y_{K})}$ acts freely on $\Xi$ by sending $(\lambda,\nu)$ to $(\lambda\mu_{M},\nu\mu)$.
@@
We first observe that the complement of $Y_{\Gamma}$ in $Y_{\Gamma}^{\ast}$ consists of finitely many points, the cusps. Furthermore, as explained in §2.2, $Y_{\Gamma}$ admits finitely many isolated singularities all of which are elliptic points. Thus there exists a smooth open $U_{K}$ of the normal projective surface $Y_{K}^{\ast}$ whose complement consists of finitely many closed points. Lemma 3.1 applied component-wise to $U_{K}$ yields $q(Y_{K}^{\ast})=\dim\operatorname{H}^{0}(U_{K},\Omega^{1}_{U_{K}})$. Let $K^{\prime}$ be any normal finite index torsion free subgroup of $K$, e.g. the intersection with the principal congruence subgroup of level $3$ (see [11, Lem. 1.4]). By Koecher’s Principle, as $Y_{K^{\prime}}{\smallsetminus}U_{K^{\prime}}$ has codimension at least $2$ in $Y_{K^{\prime}}$, we have
@@
where $U_{K^{\prime}}$ is the inverse image of $U_{K}$ under the natural projection $Y_{K^{\prime}}\to Y_{K}$. Taking invariants by the finite group $K/K^{\prime}$ in Rogawski’s formula [11, (14)] for $q(Y_{K^{\prime}}^{\ast})=\dim\operatorname{H}^{0}(Y_{K^{\prime}},\Omega^{1}_{Y_{K^{\prime}}})$ yields
$\dim\operatorname{H}^{0}(U_{K},\Omega^{1}_{U_{K}})=\sum_{(\lambda,\nu)\in\Xi}\sum_{\pi_{f}\in\Pi(\lambda_{f},\nu_{f})}\dim(\pi_{f}^{K})\frac{1+W(\lambda\nu_{M}^{-1})(-1)^{s(\pi_{f})}}{2}$
(there is a misprint in loc. cit. where one should read $(1+W(\lambda\nu_{M}^{-1})(-1)^{s(\pi_{f})})$ instead of $(W(\lambda\nu_{M}^{-1})+(-1)^{s(\pi_{f})})$; it also uses the inverse notation for the base change $\nu_{M}$). The presence of this root number translates the fact that for $\pi_{f}\in\Pi(\lambda_{f},\nu_{f})$ and $\pi_{\infty}$ the unique non-tempered holomorphic representation in the local Arthur packet $\Pi(\lambda_{\infty},\nu_{\infty})$, $\pi=\pi_{f}\otimes\pi_{\infty}$ is automorphic if, and only if, $W(\lambda\nu_{M}^{-1})=(-1)^{s(\pi_{f})}$. Both $\dim(\pi_{f}^{K})$ and $1+W(\lambda\nu_{M}^{-1})(-1)^{s(\pi_{f})}$ being preserved by the action of $\widehat{\pi_{0}(Y_{K})}$, one deduces the desired formula for $q(Y_{\Gamma}^{\ast})$ as in [11, (15)]. ∎
@@ 3.3. Twists of canonical characters and root numbers
Hecke characters $(\lambda,\nu)\in\Xi$ whose local components at each finite place have ‘minimal’ ramification are intimately related to the canonical characters studied by Gross and Rohrlich. They play a pivotal role in our production of automorphic forms contributing to the irregularity of the Picard modular surfaces of low level. We will now briefly recall some of their properties under the running assumption that $D>3$ is odd. Consider the character $\lambda_{\infty}(z)=\bar{z}\cdot|z|^{-1}$ of $M_{\infty}^{\times}\simeq\mathbf{C}^{\times}$ and let $\lambda_{f}:\widehat{\mathcal{O}}_{M}^{\times}\to\mathbf{C}^{\times}$ be a continuous character whose restriction to $\mathcal{O}_{M,p}^{\times}$ is given by the unique quadratic character
@@
for all $p$ dividing $D$, and is trivial otherwise. As $\left(\tfrac{-1}{D}\right)=-1$, it follows that $\lambda_{\infty}$ and $\lambda_{f}$ agree on $\mathcal{O}_{M}^{\times}=\{\pm 1\}$. The finiteness of the ideal class group $\mathscr{C}\!\ell_{M}=\mathbf{A}_{M}^{\times}/M^{\times}\widehat{\mathcal{O}}_{M}^{\times}M_{\infty}^{\times}$ guarantees that the resulting character of $M^{\times}\widehat{\mathcal{O}}_{M}^{\times}M_{\infty}^{\times}$ extends to a character $\lambda$ of $\mathbf{A}_{M}^{\times}$ and clearly two such extensions must differ by an ideal class character. As by construction the restriction of $\lambda_{f}$ to $\widehat{\mathbf{Z}}^{\times}$ agrees with the quadratic Dirichlet character $\left(\tfrac{\cdot}{D}\right)$ viewed as a character of $\mathbf{A}^{\times}/\mathbf{Q}^{\times}\mathrm{Nm}(\mathbf{A}_{M}^{\times})=\operatorname{Gal}(M/\mathbf{Q})$ and $\mathbf{A}^{\times}=\mathbf{Q}^{\times}\widehat{\mathbf{Z}}^{\times}\mathbf{R}_{+}^{\times}$, it follows that the restriction of $\lambda$ to $\mathbf{A}^{\times}$ equals $\left(\tfrac{\cdot}{D}\right)$, i.e., $\lambda$ is conjugate-symplectic. Such a character is called canonical and we will denote it by $\lambda_{c}$, remembering that it is only unique up to a multiplication by a character of $\mathscr{C}\!\ell_{M}$. The root number $W(\lambda_{c}^{3})=-W(\lambda_{c})=\left(\tfrac{-2}{D}\right)$ is $1$ if, and only if, $D\equiv 3\pmod{8}$ (see [28]).
@@
Assume henceforth that $\det(K)=\widehat{\mathcal{O}}_{M}^{1}$, so that $\pi_{0}(Y_{K})=\mathscr{C}\!\ell_{M}^{1}:=\mathbf{A}_{M}^{1}/M^{1}\widehat{\mathcal{O}}_{M}^{1}M_{\infty}^{1}$, and that $\widehat{\mathcal{O}}_{M}^{1}$ embeds centrally into $K$, the central character $\omega=\nu\cdot\lambda_{|M^{1}}$ of any $\pi$ contributing to (17) has to be everywhere unramified, i.e.,
@@
(18) $q(Y_{\Gamma}^{\ast})=\frac{1}{|\mathscr{C}\!\ell_{M}^{1}|}\sum_{\chi\in\Xi^{1},\omega\in\widehat{\mathscr{C}\!\ell_{M}^{1}}}\sum_{\pi_{f}\in\Pi_{f}\left(\lambda_{c}\chi_{M},{\lambda^{-1}_{c}}_{\!\!\!\!\!|M^{1}}\chi^{-2}\omega\right)}\dim(\pi_{f}^{K})\frac{1+W(\lambda_{c}^{3}\chi^{3}_{M})(-1)^{s(\pi_{f})}}{2},$
@@
where $\Xi^{1}$ denotes the set of finite order characters of $\mathbf{A}_{M}^{1}/M^{1}$ (see [28, (3)] for the fact that multiplication by an ideal class character does not change the root number).
@@
If $3$ does not divide the class number $h:=|\mathscr{C}\!\ell_{M}|$, then the action of $\mu\in\widehat{\mathscr{C}\!\ell_{M}^{1}}$ sending $(\chi,\omega)$ to $(\chi\mu,\omega\mu^{3})$ allows one to twist out the central character $\omega$ and obtain the simpler formula
(19) $q(Y_{\Gamma}^{\ast})=\sum_{\chi\in\Xi^{1}}\sum_{\pi_{f}\in\Pi_{f}(\lambda_{c}\chi_{M})}\dim(\pi_{f}^{K})\frac{1+W(\lambda_{c}^{3}\chi^{3}_{M})(-1)^{s(\pi_{f})}}{2}.$
@@
Successfully applying (18) requires one to understand how root numbers behave under twisting. As we are interested in creating irregularity at level $\Gamma^{\prime\prime}_{0}(\ell^{r})$, we focus on characters $\chi$ which are only ramified at the fixed inert prime $\ell$.
@@
For $\chi\in\Xi^{1}$ of Artin conductor $\ell^{a(\chi)}$ and $\lambda$ an odd power of a canonical character,
$W(\lambda\chi_{M})=(-1)^{a(\chi)}\chi_{\ell}(-1)W(\lambda).$
@@
Using the factorization of root numbers $W=\prod\limits_{v}W_{v}$, it suffices to prove that
@@
(21) $W_{v}(\lambda_{v}\chi_{M,v})$ $=W_{v}(\lambda_{v}),\,\text{ for all } v\neq\ell,$
where the local factors are defined using the standard additive character $\psi_{M}=\psi_{\mathbf{Q}}\circ\operatorname{Tr}_{M/\mathbf{Q}}$. Applying [29, Prop. 3] to both $\lambda_{\ell}$ and $\lambda_{\ell}\chi_{M,\ell}$ yields (20). As $\chi_{M,\infty}=1$, it suffices to check (21) for $v$ finite. Moreover, the characters $\lambda_{v}$ and $\chi_{M,v}$ are unramified for $v\nmid\ell D$, hence both sides of (21) are $1$. Finally, for $v$ dividing $D$, $\chi_{M,v}$ is unramified while $\lambda_{v}$ is tamely ramified and $\psi_{M_{v}}$ has conductor $1$, implying by [7, (5.5.1)] that $W_{v}(\lambda_{v}\chi_{M,v})/W_{v}(\lambda_{v})=\chi_{M,v}(\xi^{1+1\cdot 1})=1$. ∎